# Inference for the stochastic FitzHugh-Nagumo model from real action potential data via approximate Bayesian computation

### v1

@@ Appendix A Impact of the proposal kernel
In this section, we first define the Gaussian olcm proposal sampler introduced in [15], and then compare its performance with the Gaussian standard proposal sampler used throughout the paper (and also proposed in [15]). Alternative sequential algorithms, such as sequential important sampling ABC (SIS-ABC), and several guided sequential ABC schemes (both SIS-ABC and SMC-ABC), recently proposed in [46], have also been considered. As their performance is comparable to that of the SMC-ABC algorithm with the olcm sampler (and better when using canonical summaries, which however we discourage to choose), we do not report them here.
A.1 Gaussian optimal local covariance matrix (olcm) sampler
@@
The Gaussian standard proposal sampler $K_{r}(\cdot|\theta_{j})=\mathcal{N}(\theta_{j},2\hat{\Sigma}_{r-1})$ considered throughout the paper is characterised by a mean $\theta_{j}$ sampled from the weighted set $\{(\theta_{r-1}^{(j)},\omega_{r-1}^{(j)})_{j=1}^{N}\}$, and a covariance matrix which is not specific for $\theta_{j}$, as it is computed by using all kept particles at the previous iteration (see Section 3.2.5). The Gaussian olcm proposal sampler instead, is a perturbation kernel with a covariance matrix specific for the sampled $\theta_{j}$. Among the $N$ accepted particles at iteration $r-1$, consider the subset of $N_{0}\leq N$ weighted particles whose corresponding distances (between the simulated and the reference data) are also smaller than $\delta_{r}$, i.e.
@@
where $\bar{\mu}_{r-1}$ is a normalisation constant such that $\sum_{l=1}^{N_{0}}\mu_{r-1,l}=1$. As we choose the new threshold level $\delta_{r}$ as the $p$th percentile of the $N$ accepted distances at iteration $r-1$ (see Section 3.2.6), we already know that the set above will contain the corresponding $N_{0}=p\%N$ particles. Then, at iteration $r$, the olcm proposal is given by $K^{\textrm{olcm}}_{r}(\cdot|\theta_{j})=\mathcal{N}(\theta_{j},\Sigma^{\mathrm{olcm}}_{\theta_{j}})$, where $\Sigma^{\mathrm{olcm}}_{\theta_{j}}=\sum_{l=1}^{N_{0}}\mu_{t-1,l}(\tilde{\theta}_{r-1,l}-\theta_{j})(\tilde{\theta}_{r-1,l}-\theta_{j})^{\top}$. Accepted particles are given unnormalized weights $\tilde{w}^{(j)}_{r}$ as in Line 24 of Algorithm 1, with $k_{r}(\cdot|\theta_{r-1}^{(l)})$ being now the density of $\mathcal{N}(\theta_{r-1}^{(l)},\Sigma^{\textrm{olcm}}_{\theta_{r-1}^{(l)}})$. As the constructed covariance matrix $\Sigma^{\mathrm{olcm}}_{\theta_{j}}$ is now specific for each particle $\theta_{j}$, it will have to be computed for every $j$th particle, $j=1,\ldots,N$, introducing thus a small but non-negligible overhead in the computations.
A.2 Inference under different proposal samplers
@@
Figure 10: Computational budget, acceptance rate, runtime (in seconds), and ESS as a function of the tolerance level $\delta$ when using the SBP SMC-ABC algorithm with either the standard (black lines) or the olcm (red lines) sampler.
@@
In this section, we investigate the impact of the Gaussian standard and olcm proposal samplers on the inference results, considering not only the obtained marginal ABC posterior densities, but also other quantities of interest. In particular, in Figure 10, we report the computational budget, acceptance rates (estimated as $N$ divided by the number of sampled particles at iteration $r$), the ESS (approximated at iteration $r$ as $1/(\sum_{j=1}^{N}(w_{r}^{(j)})^{2})$), and the algorithm runtimes as a function of the tolerance level $\delta$. Remember that the first iteration of the algorithm consists on acceptance-rejection ABC, so the perturbation kernel does not play a role there. Moreover, as the threshold levels $\delta$ are automatically decreased across iterations (with $\delta_{1}\approx 0.35$), the algorithm is “evolving” (in terms of successive iterations) towards smaller values of $\delta$, so from right to left in the plots of Figure 10. By looking at the first panel, we can see that olcm needs fewer model simulations than standard to achieve a certain tolerance level, which is consistent with the higher acceptance rates (second panel) across all algorithm iterations, as soon as the proposal samplers are used, i.e. from the second iteration. For example, olcm and standard achieve an approximate threshold level of 0.025 and 0.075, respectively, for a computational budget of $10^{5}$ model simulations, requiring approximately 296000 and 888000 model simulations to achieve a threshold around 0.02. This is also consistent with the lower runtimes (obtained when running the algorithms on a desktop machine Intel Core i7-9700 CPU 3GHz RAM 32 GB) of olcm than standard (third panel). When looking at the ESS (fourth panel), a number in $[1,N]$ measuring the effectiveness of the sampler, i.e., how many particles are “relevant” at each iteration, so the higher the better, we can see that, while high for both perturbation kernels, standard has higher ESS than olcm throughout the entire run of the algorithm.

### v2

@@ Appendix C Impact of the proposal kernel
In this section, we first define the Gaussian olcm proposal sampler introduced in [17] and then compare its performance with the Gaussian standard proposal sampler used throughout the paper (and also proposed in [17]). Alternative sequential algorithms, such as SIS-ABC and several guided sequential ABC schemes (both SIS-ABC and SMC-ABC), recently proposed in [49], have also been considered. As their performance is comparable to that of the SMC-ABC algorithm with the olcm sampler (and better when using canonical summaries, which however we discourage to choose), we do not report them here.
C.1 Gaussian optimal local covariance matrix (olcm) sampler
@@
The Gaussian standard proposal sampler $K_{r}(\cdot|\theta_{j})=\mathcal{N}(\theta_{j},2\hat{\Sigma}_{r-1})$ considered throughout the paper is characterised by a mean $\theta_{j}$ sampled from the weighted set $\{(\theta_{r-1}^{(j)},\omega_{r-1}^{(j)})_{j=1}^{N}\}$ and a covariance matrix which is not specific for $\theta_{j}$, as it is computed by using all kept particles at the previous iteration (see Section 3.2.5). The Gaussian olcm proposal sampler instead, is a perturbation kernel with a covariance matrix specific for the sampled $\theta_{j}$. Among the $N$ accepted particles at iteration $r-1$, consider the subset of $N_{0}\leq N$ weighted particles whose corresponding distances (between the simulated and the reference data) are also smaller than $\delta_{r}$, i.e.
@@
where $\bar{\mu}_{r-1}$ is a normalisation constant such that $\sum_{l=1}^{N_{0}}\mu_{r-1,l}=1$. As we choose the new threshold level $\delta_{r}$ as the $p$th percentile of the $N$ accepted distances at iteration $r-1$ (see Section 3.2.6), we already know that the set above will contain the corresponding $N_{0}=p\%N$ particles. Then, at iteration $r$, the olcm proposal is given by $K^{\textrm{olcm}}_{r}(\cdot|\theta_{j})=\mathcal{N}(\theta_{j},\Sigma^{\mathrm{olcm}}_{\theta_{j}})$, where $\Sigma^{\mathrm{olcm}}_{\theta_{j}}=\sum_{l=1}^{N_{0}}\mu_{t-1,l}(\tilde{\theta}_{r-1,l}-\theta_{j})(\tilde{\theta}_{r-1,l}-\theta_{j})^{\top}$. Accepted particles are given unnormalized weights $\tilde{w}^{(j)}_{r}$ as in Line 25 of Algorithm 1, with $k_{r}(\cdot|\theta_{r-1}^{(l)})$ being now the density of $\mathcal{N}(\theta_{r-1}^{(l)},\Sigma^{\textrm{olcm}}_{\theta_{r-1}^{(l)}})$. As the constructed covariance matrix $\Sigma^{\mathrm{olcm}}_{\theta_{j}}$ is now specific for each particle $\theta_{j}$, it will have to be computed for every $j$th particle, $j=1,\ldots,N$, introducing thus a small but non-negligible overhead in the computations.
C.2 Inference under different proposal samplers
@@
Figure 14: Computational budget, acceptance rate, runtime (in seconds) and ESS as a function of the tolerance level $\delta$ when using the SBP SMC-ABC algorithm with either the standard (black lines) or the olcm (red lines) sampler.
@@
In this section, we investigate the impact of the Gaussian standard and olcm proposal samplers on the inference results, considering not only the obtained marginal ABC posterior densities, but also other quantities of interest. In particular, in Figure 14, we report the computational budget, acceptance rates (estimated as $N$ divided by the number of sampled particles at iteration $r$), the ESS (approximated at iteration $r$ as $1/(\sum_{j=1}^{N}(w_{r}^{(j)})^{2})$) and the algorithm runtimes as a function of the tolerance level $\delta$. Remember that the first iteration of the algorithm consists of acceptance-rejection ABC, so the perturbation kernel does not play a role there. Moreover, as the threshold levels $\delta$ are automatically decreased across iterations (with $\delta_{1}\approx 0.35$), the algorithm is “evolving” (in terms of successive iterations) towards smaller values of $\delta$, so from right to left in the plots of Figure 14. By looking at the first panel, we can see that olcm needs fewer model simulations than standard to achieve a certain tolerance level, which is consistent with the higher acceptance rates (second panel) across all algorithm iterations, as soon as the proposal samplers are used, i.e. from the second iteration. For example, olcm and standard achieve an approximate threshold level of 0.024 and 0.047, respectively, for a computational budget of $10^{5}$ model simulations, requiring approximately 390000 and 1562000 model simulations to achieve a threshold around 0.02 in 100 $s$ and 256 $s$, respectively. This is also consistent with the lower runtimes of olcm than standard (third panel). When looking at the ESS (fourth panel), a number in $[1,N]$ measuring the effectiveness of the sampler, i.e., how many particles are “relevant” at each iteration, so the higher the better, we can see that, while high for both perturbation kernels, standard has higher ESS than olcm throughout the entire run of the algorithm.
@@
Appendix D Real data study: ABC posterior scatter plots
In this section, we provide a more comprehensive picture of the ABC posterior distributions obtained from the four investigated action potentials 1553, 1608 (animal at rest) and 1554, 1609 (animal stimulated), cf. Section 5.
Figure 17 reports scatter plots of pairs of the kept ABC posterior samples, the black and red dots referring to the “resting” and “stimulating” scenarios, respectively. The posterior point clouds of the two “resting” (resp. “stimulating”) scenarios are located at similar parameter regions. Those regions differ across the two scenarios, with the kept ABC posterior samples in the “stimulating” case shifted towards larger parameter values. This is in agreement with what observed when looking at the marginal posterior densities reported in Figure 8 of Section 5.
Table 5 reports the corresponding correlation coefficients among pairs of kept ABC posterior samples. All four action potential datasets yield ABC posterior samples with a strong correlation among the parameters $\gamma$ and $\beta$, $\gamma$ and $\sigma$, and $\beta$ and $\sigma$. The correlation among $\gamma$ and $\sigma$ and among $\beta$ and $\sigma$ is slightly larger for the two “stimulating” scenarios compared to the “resting” scenarios. In addition, we also observe differences in the pairwise parameter correlations within the two settings. For example, we obtain a rather strong correlation between $\epsilon$ and $\gamma$ for the “resting” scenario 1553, and a rather weak one for the “resting” scenario 1608. Similarly, we observe a rather strong (resp. weak) correlation between $\epsilon$ and $\gamma$ for the “stimulating” scenario 1609 (resp. 1554).
Figure 17: Scatter plots of pairs of the kept ABC posterior samples obtained from the real action potential datasets 1553 (animal at rest, black dots in the top panels), 1554 (animal stimulated, red dots in the top panels), 1608 (animal at rest, black dots in the bottom panels) and 1609 (animal stimulated, red dots in the bottom panels).
Table 5: ABC posterior correlation coefficients of pairs of parameters, obtained under four different real action potential datasets.
case ABC posterior correlations
Rest. 1553 $\epsilon,\gamma$: 0.755; $\epsilon,\beta$: 0.668; $\epsilon,\sigma$: 0.507; $\gamma,\beta$: 0.977; $\gamma,\sigma$: 0.828; $\beta,\sigma$: 0.923
Rest. 1608 $\epsilon,\gamma$: 0.330; $\epsilon,\beta$: 0.326; $\epsilon,\sigma$: 0.443; $\gamma,\beta$: 0.985; $\gamma,\sigma$: 0.715; $\beta,\sigma$: 0.816
Stim. 1554 $\epsilon,\gamma$: 0.430; $\epsilon,\beta$: 0.417; $\epsilon,\sigma$: 0.522; $\gamma,\beta$: 0.986; $\gamma,\sigma$: 0.887; $\beta,\sigma$: 0.943
Stim. 1609 $\epsilon,\gamma$: 0.750; $\epsilon,\beta$: 0.734; $\epsilon,\sigma$: 0.760; $\gamma,\beta$: 0.985; $\gamma,\sigma$: 0.908; $\beta,\sigma$: 0.958